OpenStudy (anonymous):

Algebra 1 Homework Help?

OpenStudy (anonymous):

I need some help figuring out math problems for my homework project, I can't afford to get this wrong since my grade is low. Help? Snapshot of my textbook here (problem 1):

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Yup? xD

OpenStudy (anonymous):

which one do u want help with first?

OpenStudy (anonymous):

The first equation in the picture. 4.

OpenStudy (anonymous):

ok, so the first thing we see is that we can factor out an x from every term

OpenStudy (anonymous):

which would give us 2x + 6 / 10x^2

OpenStudy (anonymous):

do u see what i did?

OpenStudy (anonymous):

Not clearly. Can you explain your steps?

OpenStudy (anonymous):

sure. notice all parts have an x right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so, one thing you can do to simplify a fraction is to divide the numerator (top number) and the denominator (bottom number) by the same value

OpenStudy (anonymous):

you can separate them since they have the same denominator like \[\frac{ x+x^3 }{ x^2 } \rightarrow \frac{ x }{ x^2 }+\frac{ x^3 }{ x^2 }\]

OpenStudy (anonymous):

i don't think you should do that

OpenStudy (anonymous):

in this, since they have the same base which is X, just subtract the exponent which is 1-2 and 3-2: the final answer would be \[x^{-1}+x\]

OpenStudy (anonymous):

there are tons of approach we can do for this problem, it's just which one a person finds it easier :)

OpenStudy (anonymous):

I don't find any of it easy. I'm an English writer, not a math wizard :c I'm still lost on how to actually work the problem fully, I actually don't know how o_o

OpenStudy (anonymous):

ok, lets divide the top part by x. that gives us 2x + 6 (x)(2x + 6) = 2x^2 + 6 now lets divide x from the bottom. that gives us 10x^2 (x)(10x^2) = 10x^3

OpenStudy (anonymous):

Oh! Okay, that makes sense actually.

OpenStudy (anonymous):

So would the second equation 5. be worked the same way?

OpenStudy (anonymous):

hold on we are not done yet. notice how all the numbers (coeficients) we have are even? we can divide them all by 2. so the final answer is (x + 3)/ 5x^2

OpenStudy (anonymous):

Oh... I didn't know that...

OpenStudy (anonymous):

its the same idea though, you look for numbers or variables that you can divide out

OpenStudy (anonymous):

So 5. would be m = 3?